OpenStudy (anonymous):

Simplify the rational expression. (4x^2 − 20x)/(x^3 − 125)

OpenStudy (anonymous):

for the numerator 4x^2 - 20x = 4x(x-5) for the denominator x^3 - 125 = x^3 - 5^3 = (x-5)(x^2 + 5x + 25) there fore 4x(x-5)/(x-5)(x^2+5x+25) = 4x/x^2+5x+25 :)

OpenStudy (anonymous):

thanks, i factored the numerator wrong